OpenStudy (anonymous):

Please help!: What are the solutions & x-intercepts? 4x^2=12 Also Give exact and approximate solutions to three decimal places. y^2-6y+9=64

OpenStudy (mertsj):

Find the x intercepts by letting y = 0. They are the roots of the equation 4x^2-12=0

OpenStudy (mertsj):

x^2-3=0 x=+-sqrt3=+-1.732

OpenStudy (mertsj):

(y-3)^2=64 y-3=+-8 y=-5,11